A box contains 5 balls. One is numbered 1, two are numbered 2, and two are numbered 3. The balls are mixed and one is selected at random. After a ball is selected, its number is recorded. Then it is replaced. If the experiment is repeated many times, find the standard deviation of the numbers on the balls


1
Expert's answer
2022-01-20T13:26:37-0500

Let X be a random variable represents the number on the ball, then

"P(X=1)=0.2"

"P(X=2)=0.4"

"P(X=3)=0.4"

"\sigma=\sqrt{D}=\sqrt{M(X^2)-M^2(X)}"

"M(X)=1*0.2+2*0.4+3*0.4=2.2\implies M^2(X)=4.84"

"M(X^2)=1^2*0.2+2^2*0.4+3^2*0.4=5.4"

So, "\sigma=\sqrt{5.4-4.84}=\sqrt{0.56}\approx0.75"